The Rioja Company (Rioja), a Spanish-based manufacturer of pleasure sailing boats, purchased a fiberglass molding machine on January 1, Year 1 for Euro 100,000. The use life of the molding machine for book purposes is 10 years on a straight-line basis with no salvage value. The Spanish tax authorities allow for depreciation for this type of machine over 5 years on a straight-line basis. The corporate tax rate in Spain is 35%. The book basis, the tax basis, and the balance in the deferred tax liability account for this machine for the first 5 years were as follows, respectively:

Book cost: Euro 100,000 each year

Book Accumulated Depreciation: Euro 10,000 at the end of Year 1, increasing by Euro 10,000 each year for Years 2-5

Book basis: Book cost less book accumulated depreciation, respectively, for each year using the above information for Years 1-5

Tax Cost: Euro 100,000 each year

Tax Accumulated Depreciation: Euro 20,000 at the end of Year 1, increasing by Euro 20,000 each year for Years 2-5

Tax basis: Tax cost less tax accumulated depreciation

Deferred tax liability: Euro 3,500 at end of Year 1; increasing by Euro 3,500 each year from Years 2-5

Required:

On December 31, Year 3, after recording its year-end journal entries, Rioja receives an offer to sell the molding machine to an Italian-based sailboat company, Vino, for Euro 120,000. Rioja has a very short window to consider what its options are for this machine, as Vino would like to complete the transaction on January 1, Year 4.

You should include any required supporting schedules and each required journal entry to help explain the income tax ramifications for Rioja’s options under both the US GAAP and IFRS. Each journal entry should be accompanied by an explanation of how you calculated the dollar amount of the respective journal entry and be supported by reference(s) to the respective IFRS and FASB literature.

Once your analysis is complete, what will be your recommendation to management of Rioja —should it accept Vino’s offer or not, and why or why not? Ensure your recommendation is specific.
Answer #1

“Applicable accounting standard is IAS 16- Property, Plant and Equipment”

After 3 Years, Machine value will be

  • As per book = 1,00,000(10,000)*3= 70,000 Euros
  • As per Tax   = 1,00,000(20,000)*3 =40,000 Euros

And deferred tax liability in books after 3 yrs = 3,500*3=10,500 Euros

Offered value from Vino’s = 1,20,000 Euros

From the above values, it is very clear that Offered price is more than both WDV’s(Written down value / Net asset value) as per Book and Tax basis. So it's better to sell the Machine as it yielding Profit under both bases.

Accounting implications under

-Book basis:    Bank account   Dr    1,20,000

                                   To Machine account 70,000

                                   To Profit on Sale a/c   50,000

-Tax Basis:   Bank account Dr 1,20,000

                           To Machine a/c                40,000

                            To Profit on sale a/c      80,000

But Tax impact on sale of Machine is same under Both basis:

  • Tax burden under Book basis = 50,000*35%+ 10,500(Def Tax Liab) =28,000
  • Tax burden under Tax a/c basis = 80,000*35%=     28,000.